OpenStudy (anonymous):

A spinner has 9 equal sections labeled 1 through 9. What is the probability of spinning an odd number on the spinner and tossing heads on a coin? A.1/18 B.5/18 C.1/3 D.6/11

OpenStudy (anonymous):

From 1-9, there are 5 odd numbers (1, 3, 5, 7, 9) Prob of odd = (5/9) Heads up = (1/2) (5/9)*(1/2) = (5/18)

OpenStudy (anonymous):

you helped a ton, thanks!